OpenStudy (fuad):

quantum number if n=5.,what are the orbitals?

OpenStudy (abhisar):



OpenStudy (abhisar):

Similarly we can go up to l=n-1 i.e. in your case l=5-1 = 4, it will give you 25 orbitals (no. of orbitals in nth shell = \(\sf n^2\)).

OpenStudy (fuad):

then what would be the sub orbit

OpenStudy (abhisar):

Do you mean Sub-Shell, like S sub shell, P sub shell?

OpenStudy (fuad):

yes

OpenStudy (abhisar):

For nth shell, possible shells are 0 to n-1, in your case they will be 0, 1, 2, 3 and 4 which corresponds to s, p, d, f and g respectively